# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ 4 Theoretical Results on CL in Overparameterized Regimes
Continual learning can be considered a special type of MTL, with an extra constraint that the tasks arrive sequentially during a continual episode. At each timestep, $t$, the model is only exposed to $\mathcal{D}_{t}$, and the goal is to continually train a model that performs well across all tasks at the end of the episode. Note that the data for past learned tasks is not accessible when the current task is learned.
The most naive continual learner can be implemented using Equation 5, meaning that the learner only adapts to the last task at hand without any constraints. Such a learner performs well on the most recent task, but its performance on previous tasks degrades, a phenomenon called catastrophic forgetting in the CL literature. The forgetting effects emerge from the fact that the model is adapted to fit the current task data, which would make the learnable parameters sub-optimal for the past learned tasks. More formally, let $w_{{\bm{v}}{t}}$ denote the weights of the continual learner at timestep $t$ after sequentially observing tasks from $1$ to $t$. Average forgetting (lower is better) reflects the amount of negative backward knowledge transfer in a continual learner, defined as $F(w_{{\bm{v}}{T}}):=\frac{1}{T-1}\sum_{t=1}^{T-1}\left[\mathcal{L}_{t}(w_{{\bm{v}}{T}})-\mathcal{L}_{t}(w_{{\bm{v}}{t}})\right]$. For simplicity, in the rest of this section, we assume that $n_{t}=n$ for all tasks.
@@
Most successful methods in deep continual learning benefit from memory replay buffer (van de Ven et al., 2022). These methods store a small portion of the training data from each task. When learning new tasks, the stored samples are used to either regularize or retrain the model. In this section, we seek to theoretically study replay-based continual learning in the context of overparameterized linear models. Specifically, assume we have access to $m_{i}$ samples for task $i$ in the memory buffer when solving task $t+1$. Let the memory size be $\bar{m}_{t}=\sum_{i=1}^{t}m_{i}$. We use the notation $\mathbf{\hat{X}}_{1t}\in\mathbb{R}^{p\times\bar{m}_{t}}$ and $\mathbf{\hat{y}}_{1t}\in\mathbb{R}^{\bar{m}_{t}}$ to denote the training data and their labels in the memory. Then, the rehearsal-based continual learning model can be formulated as:
$w_{{\bm{v}}{{t+1}}}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad\mathbf{X}_{t+1}^{T}w=\mathbf{y}_{t+1},\quad\mathbf{\hat{X}}_{1t}^{T}w=\mathbf{\hat{y}}_{1t}.$ (12)
Replay+Regularization methods.
Another family of CL methods are regularization techniques that aim to mitigate catastrophic forgetting by imposing constraints during the learning process. These constraints typically involve modifying the learning objective by penalizing the drift in the parameters space to preserve previously acquired knowledge. Previously, regularization-based continual learning has been studied in linear models (Lin et al., 2023). However, pure regularization is often not sufficient on its own for optimal continual learning. In practice, it is combined with memory replay methods, which can be formulated as:
$w_{{\bm{v}}{{t+1}}}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w-w_{{\bm{v}}{{t}}}\|_{2}^{2}\quad\text{s.t.}\quad\mathbf{X}_{t+1}^{T}w=\mathbf{y}_{t+1},\quad\mathbf{\hat{X}}_{1t}^{T}w=\mathbf{\hat{y}}_{1t},$ (13)
where $w_{{\bm{v}}{{0}}}=\mathbf{0}\in\mathbb{R}^{p}$.
@@ 4.1 Theoretical Results on CL
In this section, we theoretically study the continual learners described in Equations 12 and 13. With a closer look, solving Equation 12 at $t=T$ is a specific case of multi-task learning described in Equation 7, achieved by setting $n_{1}=m_{1},...,n_{T-1}=m_{T-1}$ and $n_{T}=n$. Therefore, we avoid repeating the theoretical results, and instead, we provide a corollary for the two-task case.
@@
The generalization error of the regularized learner is more complicated compared to the pure replay-based model and requires more delicate investigation. In fact, the dependence of $w_{{\bm{v}}{t}}$ to the samples in the memory buffer is the source of such complication and produces many cross-terms in the final expression. To keep the results understandable and intuitive, we present the results for the two tasks case here and leave the full form for Appendix C.4. Appendix D also offers an in-depth discussion dedicated to the effect of regularization.
@@
Equations 14 to 17 reveal the effect of memory capacity. Setting $m=0$ yields the naive (regularized) sequential learner. Increasing the memory size reduces the forgetting and the generalization error through the terms $G_{2}$ and $F_{2}$, respectively. However, a larger memory size can be harmful through the terms $G_{1}$ and $F_{1}$ if the tasks are highly dissimilar. Additionally, the full MTL model is recovered when $m=n$, meaning that all samples from the previous tasks are stored in the buffer.
An interesting observation is the effect of model size on the forgetting and the generalization error. With a larger $p$, both the positive and negative terms in the forgetting vanish. This suggests that bigger models with more capacity are less vulnerable to forgetting (Goldfarb & Hand, 2023). The next observation is that a large $p$ reduces the effect of memory size. This means a natural trade-off exists in practical applications with limited physical memory. In other words, one may consider investing the hardware in deploying larger models or consider a larger memory buffer for training samples. In fact, this is an essential result that sheds light on several state-of-the-art continual learning models where they reported that sample memory is not enough and the best solution is to also keep some parts of the architecture in the memory (Zhou et al., 2023; Wang et al., 2022; Douillard et al., 2022). We revisit this result in Section 5 when studying DNNs.
Although we presented the $T=2$ versions of the theorems for better comprehension, the mentioned phenomena are also observable in the $T>2$ case as presented in Appendices C.3 and C.4. Figures 1(c) and 1(d) visualize the error behavior of the replay-based model for different values of $m$ when learning $T=10$ tasks. As demonstrated, the location of the error peak and its strength are affected by $m$. Moreover, it is surprisingly observable that the continual learner can sometimes outperform the MTL baseline even when tasks are collaborative, especially at the multi-task learner’s interpolation threshold.
@@ 5 Empirical Exploration for Deep Neural Networks
Although fully understanding studying DNNs through theory seems infeasible, recent literature has pointed out connections between DNNs optimized by SGD and linear models and used these models as a proxy to study DNNs (Please refer to Appendix E for a discussion on the connection between overparameterized DNNs and linear models). We follow a similar approach and empirically demonstrate how our results can provide insights for understanding and designing deep MTL and CL models.
@@ 5.1 Experimental Setup
Our experiments are performed using three datasets and three backbone architectures to offer the generalizability and scalability of our results to non-linear models. It is notable that all experiments are repeated at least 3 times, and the average values are reported. In the subsequent paragraphs, we present the most important details of our experiments and leave further information for Appendix F.
@@
We use CIFAR-100, ImageNet-R (IN-R), and CUB-200 datasets in our experiments. We generate MTL or CL tasks by randomly splitting these datasets into 10 tasks, with an equal number of classes in each task(van de Ven et al., 2022). No specific data augmentation or preprocessing techniques were applied, except for resizing to accommodate the input requirements of the different backbones.
@@
Our main focus is on the models operating in overparameterized regimes. Therefore, we leverage models that are complex enough to ensure that the training data is completely overfitted by the model. When using the pretrained backbones, we implement a shared 3-layer MLP and a classification head on top of the feature extractor and use the width of the MLP to manipulate the model’s complexity by changing the number of hidden neurons for all of the 3 layers. On the other hand, when the backbone is trainable, we adjust the model’s complexity by considering different values for the number of convolutional filters in the backbone. In particular, the ResNet-18 has layers with $[k,2k,4k,8k]$ number of filters and we control the width of the backbone by changing the value of $k$. The typical ResNet-18 width has $k=64$ filters. On top of this backbone, we use an MLP layer with a single hidden layer of size 128, followed by a classification head.
@@ 5.2 Empirical Results
Figure 2: Performance comparison of the single-task vs multi-task learners on CIFAR-100 using pretrained ResNet-18 backbone. The vertical axis is the cross-entropy loss, and the horizontal axis is $\log_{2}(k)$ where $k$ is the width of the MLP on top of the backbone. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels. (a) and (b) correspond to the train and test loss with the single-head classifier, while (c) and (d) show the train and test loss of the multi-head architecture.
@@
Figure 3: Test loss comparison of the single-task vs multi-task learners on different datasets with pretrained backbones. The vertical axis is the cross-entropy loss, and the horizontal axis is $\log_{2}(k)$ where $k$ is the width of the MLP on top of the backbone. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels. The multi-head architecture was used in all MTL experiments.
Multi-task vs single-task learning.
Figure 2 presents the train and test behavior of a multi-task DNN compared to single-task models, when trained on CIFAR-100 using the pretrained ResNet18. As depicted in Figures 2(a) and 2(b), train error is almost zero for heavily overparameterized models. However, the test error behavior is in line with our theoretical analysis and is similar to the linear models illustrated in Figure 1 from several aspects: (i) The training interpolation threshold and the corresponding test loss peak for STL and MTL models are observable at different locations, which is consistent with the observation 1. (ii) The error peak exists even with zero amount of label noise but is intensified with larger noise, which aligns with the observation 4. (iii) A descent floor exists in the overparameterized regime where the test loss is minimized, which is related to the observation 3. These observations are crucial when designing MTL models and should be considered when optimizing the model size.
In Figures 2(a) and 2(b), while the single-head architecture is used and although the training error is zero for large enough models, we observe that single-task learners significantly outperform multi-task learners. Inspired by the observations 2 and 4, this behavior suggests that weight sharing is not happening effectively, and the tasks are not fully collaborative. To dig deeper into this issue which can be considered a type of negative transfer learning, we analyze the optimal weight for each task per MLP layer. In other words, we start from a fully shared model, repeatedly pick and unshare the weights of a specific layer and optimize it to find the optimal weights of that layer per each task. Inspired by Equation 11, we calculate the inner product of the task-specific weights at each layer across different tasks and report the average values in Figure 4. For further details and discussion about this experiment, please refer to Appendix G.3.
Figure 4: The average across-task inner product of task-specific optimal weights at different MLP layers for different values of $k$.
According to Figure 4, we observe that the weights at the final classification layer have the most negative inner product values. This observation suggests that CIFAR-100 classes are highly conflicting in the classification layer. Nevertheless, extending the classification head to the multi-head architecture enables effective knowledge transfer in addition to an improved performance for the multi-task learner, as illustrated in Figure 2(d). In other words, while unsharing any layer can benefit the MTL performance (as elaborated in Appendix G.3), unsharing the classification head is the most effective approach, which leads to a desired behavior closely similar to the collaborative version of the linear models in Figure 1(a). As an additional experiment, we present Figure 5, which provides a similar comparison using an unfrozen version of ResNet-18 training from random initialization. Although this backbone is much deeper compared to the linear models we studied in Section3.2, the consistent behavior suggests that our findings also generalize to more complex architectures.
Figure 5: Test loss comparison of the single-task (top) vs multi-task (bottom) learners on CIFAR-100 with unfrozen ResNet-18. The vertical axis is the cross-entropy loss and the horizontal axis is $\log_{10}(k)$ where $k$ controls the number of filters in convolutional layers.
@@
Additionally, Figure 3 offers further experiments on CUB and IN-R datasets trained on two different pretrained backbones. Observing similar phenomena to what we presented in observations 1-4, endorses our hypothesis regarding the connection between our results for MTL using linear models and MTL using DNNs. For example, in all of these experiments the test error peak of the multi-task learner is considerably strong, while in some configurations, the STL error peak is negligible. This observation emphasizes that the label noise is not the sole source of the error peak; rather, as predicted from our theoretical analysis, the inherent dissimilarity across different tasks will also contribute to the strength of the peak. We conclude that our results can be used to explain the behavior of DNNs and combined with empirical investigation, they can be used for optimal design. Some complementary experiments, as well as the full training and testing curves of MTL experiments, are provided in Appendix G.1.
@@
Figure 6 presents the test loss and the forgetting of a DNN continual learner on CIFAR-100 for different amounts of memory budget and model size. These models follow a trend very similar to the linear models presented in Section 4 and Figure 1(c). In particular, there exists a test error peak for all models, but its location and strength are controlled by the memory size. By increasing the memory size, the test error peak naturally shifts to the right and its intensity is reduced. Another observation aligning with the linear models is that the forgetting greatly reduces for large model sizes, even without any memory, a phenomenon also confirmed by recent studies on regularization-based CL methods (Goldfarb & Hand, 2023).
@@
Figure 6: The test loss and forgetting of a continual learner on CIFAR-100 w.r.t. width of the model, $k$, for different levels of per-class memory capacity. $m=0$ means naive fine-tuning, and $m=500$ corresponds to the full MTL setup. (a) and (b) correspond to the experiments with the Pretrained ResNet-18 and subfigures (c) and (d) show the results when fine-tuning a ResNet-18 model from scratch. The multi-head architecture with no label noise was used in these experiments.
Additionally, it is also noticeable that increasing the memory capacity does not always uniformly increase the performance for all values of $p$. Especially, the memory size is less effective with larger models. This raises a fundamental question when designing CL models in practice. Given a fixed budget for physical memory, is it more efficient to spend it on increasing the model size or building a bigger sample memory? Table 1 contains the results of experiments with varying budgets dedicated to the replay memory buffer. As the results show, different optimal balance points exist for all datasets that maximize the average accuracy and minimize forgetting. This observation aligns with the recent CL methods, where an equal portion of the sample memory budget is devoted to specialized blocks (Zhou et al., 2023; Wang et al., 2022) or task-specific tokens (Douillard et al., 2022). Therefore, we propose that our theoretical findings in linear models can be applied to practical CL algorithms to gain a deeper understanding.
Table 1: The effect of varying sample memory budget and architecture memory budget over the average accuracy and average forgetting of overparameterized continual learners on two datasets. All values in this table are presented as percentages. The sample budget is linearly proportional to the number of training images stored in the memory, where 100% corresponds to storing 10% of the whole training set. (A complete version of this table with more details is presented in Appendix F.)
Mem. Budget Arch. Budget CIFAR-100 Imagenet-R CUB PMNIST
Avg.
Acc.
Avg.
Forg.
Avg.
Acc.
Avg.
Forg.
Avg.
Acc.
Avg.
Forg.
Avg.
Acc.
Avg.
Forg.
100 <1 $56.6_{\pm 0.4}$ $-22.5_{\pm 0.5}$ $76.1_{\pm 0.4}$ $-3.3_{\pm 0.6}$ $87.6_{\pm 0.2}$ $-4.5_{\pm 0.2}$ $32.8_{\pm 2.8}$ $-69.8_{\pm 3.0}$
80 20 $70.6_{\pm 0.1}$ $-7.1_{\pm 0.1}$ $76.9_{\pm 0.6}$ $-2.0_{\pm 0.3}$ $84.1_{\pm 0.5}$ $-6.7_{\pm 0.0}$ $82.9_{\pm 1.0}$ $-13.9_{\pm 1.1}$
60 40 $\bm{71.9_{\pm 0.2}}$ $\bm{-5.4_{\pm 0.3}}$ $77.1_{\pm 0.6}$ $-1.4_{\pm 0.4}$ $88.5_{\pm 0.3}$ $-1.8_{\pm 0.6}$ $87.7_{\pm 0.4}$ $-8.5_{\pm 0.4}$
40 60 $66.3_{\pm 0.4}$ $-11.3_{\pm 0.4}$ $\bm{78.9_{\pm 0.5}}$ $-1.0_{\pm 0.1}$ $\bm{92.0_{\pm 0.1}}$ $\bm{-0.1_{\pm 0.2}}$ $\bm{88.8_{\pm 0.1}}$ $\bm{-7.1_{\pm 0.1}}$
20 80 $67.4_{\pm 0.3}$ $-9.7_{\pm 0.3}$ $78.5_{\pm 0.3}$ $\bm{-0.6_{\pm 0.1}}$ $91.7_{\pm 0.4}$ $-0.4_{\pm 0.3}$ $88.2_{\pm 0.3}$ $-7.2_{\pm 0.2}$
0 100 $63.8_{\pm 0.3}$ $-10.8_{\pm 0.3}$ $49.5_{\pm 0.6}$ $-20.5_{\pm 0.5}$ $75.8_{\pm 0.7}$ $-12.2_{\pm 0.7}$ $68.2_{\pm 1.0}$ $-17.9_{\pm 0.2}$

### v2

@@ 4 Theoretical Results on CL in Overparameterized Regimes
Continual learning can be considered a special type of MTL, with an extra constraint that the tasks arrive sequentially during a continual episode. At each timestep $t$ the model is only exposed to $\mathcal{D}_{t}$, and the goal is to continually train a model that performs well across all tasks at the end of the episode. Note that the full data for past learned tasks is not accessible when the current task is being learned. In this case, sequentially fine-tuning on the most recent task at hand results in a model that performs well on the most recent task, but with degraded performance on previous tasks, a phenomenon which is called catastrophic forgetting. Formally, let $w_{{\bm{v}}{t}}$ denote the weights of the continual learner at timestep $t$ after sequentially observing tasks from $1$ to $t$. Average forgetting (lower is better) reflects the amount of negative backward knowledge transfer in a continual learner, defined as $F(w_{{\bm{v}}{T}}):=\frac{1}{T-1}\sum_{t=1}^{T-1}\left[\mathcal{L}_{t}(w_{{\bm{v}}{T}})-\mathcal{L}_{t}(w_{{\bm{v}}{t}})\right]$. For simplicity, in the rest of this section, we assume that $n_{t}=n$ for all tasks.
Implicit Regularization for Continual Learning.
Prior to our work, recent theoretical CL studies relied on the implicit regularization of SGD to achieve a continual learning algorithm and analyzed its characteristics (Evron et al., 2022; Lin et al., 2023; Goldfarb & Hand, 2023; Goldfarb et al., 2024). In other words, they start from a zero initialization of weights, i.e. $w_{{\bm{v}}{{0}}}=0$, and simply use SGD to fit the training data of the new tasks as the tasks arrive sequentially. Because of the implicit bias of SGD, this method is equivalent to the following optimization:
$w_{{\bm{v}}{{t+1}}}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w-w_{{\bm{v}}{{t}}}\|_{2}^{2}\quad\text{s.t.}\quad X_{t+1}^{\top}w=y_{t+1},$ (14)
where $w_{{\bm{v}}{{0}}}=0\in\mathbb{R}^{p}$. This continual learner, often called sequential finetuning in CL, is not effective enough in practice and yields high forgetting. Even further, using explicit regularization can not solve the issue in practice and therefore, there is a need to study more useful CL strategies.
@@
Most successful methods in deep continual learning benefit from memory replay buffer (van de Ven et al., 2022). These methods store a small portion of the training data from each task. When learning new tasks, the stored samples are used to either regularize or retrain the model. In this section, we seek to theoretically study replay-based continual learning in the context of overparameterized linear models. Specifically, assume we have access to $m_{i}$ samples for task $i$ in the memory buffer when solving task $t+1$. Let the memory size be $\bar{m}_{t}=\sum_{i=1}^{t}m_{i}$. We use the notation $\hat{X}_{1:t}\in\mathbb{R}^{p\times\bar{m}_{t}}$ and $\hat{y}_{1:t}\in\mathbb{R}^{\bar{m}_{t}}$ to denote the training data and their labels in the memory.
A simple strategy to update the weights, in this case, is to ignore prior weights learned so far, start again from the zero-initialized weights, and use SGD to fit both the data of the new task as well as the older data stored in the memory buffer. The described replay-based continual learning optimization is equivalent to
$w_{{\bm{v}}{{t+1}}}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w\|_{2}^{2}\quad\text{s.t.}\quad X_{t+1}^{\top}w=y_{t+1},\quad\hat{X}_{1:t}^{\top}w=\hat{y}_{1:t}.$ (15)
As opposed to Equation 14, the continual learner introduced here, tends toward low-norm solutions, but does not seek to minimize its distance with the weights acquired so far from previous tasks. Therefore, we name this strategy as pure replay-based continual learning.
Implicit Regularization+Replay-Based Continual Learning.
Lastly, we can think of a continual learner that uses a mixed strategy of relying on both the replay memory buffer as well as the implicit regularization, i.e., to store older samples in the memory and also sequentially update the weights using SGD without reinitializing to zero. This strategy can be formulated as solving
$w_{{\bm{v}}{{t+1}}}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w-w_{{\bm{v}}{{t}}}\|_{2}^{2}\quad\text{s.t.}\quad X_{t+1}^{\top}w=y_{t+1},\quad\hat{X}_{1:t}^{\top}w=\hat{y}_{1:t},$ (16)
where $w_{{\bm{v}}{{0}}}=0\in\mathbb{R}^{p}$. Notice that the optimization described in Equation 14 is a special case of Equation 16 when $m=0$.
@@ 4.1 Theoretical Results on CL
In this section, we theoretically study the continual learners described in Equations 15 and 16. With a closer look, solving Equation 15 at $t=T$ is a specific case of multi-task learning described in Equation 9, achieved by setting $n_{1}=m_{1},...,n_{T-1}=m_{T-1}$ and $n_{T}=n$. Therefore, we avoid repeating the theoretical results, and instead, we provide a corollary for the two-task case.
@@
The generalization error of the regularization+replay-based learner is more complicated compared to the pure replay-based model and requires more delicate investigation. In fact, the dependence of $w_{{\bm{v}}{t}}$ to the samples in the memory buffer is the source of such complication and produces many cross-terms in the final expression. To keep the results understandable and intuitive, we present the results for the two tasks case here and leave the full form for Appendix A.4.
@@
Equations 17 to 20 reveal the effect of memory capacity. Setting $m=0$ yields the results for the sequential learner described in Equation 14. Increasing the memory size reduces the forgetting and the generalization error through the terms $G_{2}$ and $F_{2}$, respectively. However, a larger memory size can be harmful through the terms $G_{1}$ and $F_{1}$ if the tasks are highly dissimilar. Additionally, the full MTL model is recovered when $m=n$, meaning that all samples from the previous tasks are stored in the buffer.
An interesting observation is the effect of model size on the forgetting and the generalization error. With a larger $p$, both the positive and negative terms in the forgetting vanish. This observation suggests that bigger models with more capacity are less vulnerable to forgetting (Goldfarb & Hand, 2023). Moreover, the next observation is that a large $p$ reduces the effect of memory size which means a natural trade-off exists in practical applications with limited physical memory. In other words, one may consider investing the hardware in deploying larger models or consider a larger memory buffer for training samples. In fact, this is an essential result that sheds light on several state-of-the-art continual learning models where they reported that sample memory is not enough and the best solution is to also keep some parts of the architecture in the memory (Zhou et al., 2023; Wang et al., 2022; Douillard et al., 2022). We revisit this result in Section 5 when studying DNNs.
Although we presented the $T=2$ versions of the theorems for better comprehension, the mentioned phenomena are also observable in the $T>2$ case as presented in Appendices A.3 and A.4. Figures 1(c) and 1(d) visualize the error behavior of the replay-based model for different values of $m$ when learning $T=10$ tasks. As demonstrated, the location of the error peak and its strength are affected by memory size, $m$. Moreover, it is surprisingly observable that the continual learner can sometimes outperform the MTL baseline even when tasks are collaborative, especially at the multi-task learner’s interpolation threshold.
@@ 5 Empirical Exploration for Deep Neural Networks
5.1 Empirical Results
Although fully understanding studying DNNs through theory seems infeasible, recent literature has pointed out connections between DNNs optimized by SGD and linear models and used these models as a proxy to study DNNs (Please refer to Appendix B for a discussion on the connection between overparameterized DNNs and linear models). We follow a similar approach and empirically demonstrate how our results can provide insights for understanding and designing deep MTL and CL models. In summary, our most important observations are as follows:
  • MTL DNNs experience a test error peak at the interpolation threshold, but unlike STL models, the peak not only depends on the label noise but is also intensified when tasks are highly dissimilar or ineffective parameter sharing is happening (Figure 2), a consistent behavior across various practical datasets and architecture (Figures 4 and 5).
  • The distance of optimal task-specific parameters correlates with the amount of conflict that is happening across different tasks and can provide evidence for deciding which layers to decouple (Figure 3).
  • Memory-based continual learners also undergo a double descent as the number of parameters is increased. The characteristics of the error peak depend on the memory size as well as task-similarity and training size (Figure 6).
  • High overparametrization can diminish the negative knowledge transfer (forgetting). Additionally, overparametrized models benefit less from explicit memory buffers. Therefore, it is important to adjust the resources accordingly between model parameters and raw training samples (Table 1).
@@ 5.2 Experimental Setup
Our experiments are performed using three datasets and three backbone architectures to offer the generalizability and scalability of our results to non-linear models. It is notable that all experiments are repeated at least 3 times, and the average values are reported. In the subsequent paragraphs, we present the most important details of our experiments and leave further information for Appendix C.
@@
We use CIFAR-100, ImageNet-R (IN-R), and CUB-200 datasets in our experiments. We generate MTL or CL tasks by randomly splitting these datasets into 10 tasks, with an equal number of classes in each task (van de Ven et al., 2022). No specific data augmentation or preprocessing techniques were applied, except for resizing to accommodate the input requirements of different backbones.
@@
Our main focus is on the models operating in overparameterized regimes. Therefore, we leverage models that are complex enough to ensure that the training data is completely overfitted. When using the pretrained backbones, we implement a shared 3-layer MLP on top of the feature extractor followed by a classification head. We use the width of the MLP to manipulate the model’s complexity by changing the number of hidden neurons for all of the 3 layers. On the other hand, when the backbone is trainable, we adjust the model’s complexity by considering different values for the number of convolutional filters in the backbone. In particular, the ResNet-18 has layers with $[k,2k,4k,8k]$ number of filters and we control the width of the backbone by changing the value of $k$. The typical ResNet-18 width has $k=64$ filters. On top of this backbone, we use an MLP layer with a single hidden layer of size 128, followed by a classification head.
@@
Figure 2: Comparing the performance of the single-task vs multi-task learners on CIFAR-100 using pretrained ResNet-18 backbone. The vertical axis is the cross-entropy loss and the horizontal axis is the width of the MLP on top of the backbone. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels. (a) and (b) correspond to the train and test loss with the single-head classifier, while (c) and (d) show the train and test loss of the multi-head architecture. Similar to linear models studied previously, deep MTL models also undergo the double descent behaviors with the test error peak depending on the characteristics of the task similarity and strength of the noise.
Figure 3: The average across-task inner product of task-specific optimal weights at different layers for different values of MLP width, $k$. As this figure shows, the weights of the classification head has the most negative inner products which means that the tasks are highly conflicting in this layer.
Multi-task vs single-task learning.
Figure 2 presents the train and test behavior of a multi-task DNN compared to single-task models, when trained on CIFAR-100 using the pretrained ResNet18. As depicted in Figures 2(a) and 2(b), train error is almost zero for heavily overparameterized models. However, the test error behavior is in line with our theoretical analysis and is similar to the linear models illustrated in Figure 1 from several aspects: (i) The training interpolation threshold and the corresponding test loss peak for STL and MTL models are observable at different locations, which is consistent with the observation 1. (ii) The error peak exists even with zero amount of label noise but is intensified with larger noise, which aligns with the observation 3. (iii) A descent floor exists in the overparameterized regime where the test loss is minimized, which is related to the observation 4. These observations are crucial when designing MTL models and should be considered when optimizing the model size.
In Figures 2(a) and 2(b), while the single-head architecture is used and although the training error is zero for large enough models, we observe that single-task learners significantly outperform multi-task learners. Inspired by the observations 2 and 3, this behavior suggests that weight sharing is not happening effectively, and the tasks are not fully collaborative, which can be considered a type of negative transfer learning. To dig deeper into this issue, we analyze the optimal weight for each task per MLP layer. In other words, we start from a fully shared model, repeatedly pick and unshare the weights of a specific layer and optimize it to find the optimal weights of that layer per each task. Inspired by Equation 13, we calculate the inner product of the task-specific weights at each layer across different tasks and report the average values in Figure 3. For further details and discussion regarding this experiment, please refer to Appendix D.3.
Figure 4: Test loss of the single-task (Top) vs multi-task (Bottom) learners on CIFAR-100 with unfrozen ResNet-18. The vertical axis is the cross-entropy loss and the horizontal axis is the scale factor that controls the number of filters in convolutional layers. As we observe here, the test loss behaves similarly to the case where only the MLP head is trained.
According to Figure 3, we observe that the weights at the final classification layer have the most negative inner product values. This observation suggests that CIFAR-100 classes are highly conflicting in the classification layer. Nevertheless, extending the classification head to the multi-head architecture enables effective knowledge transfer in addition to an improved performance for the multi-task learner, as illustrated in Figure 2(d). In other words, while unsharing any layer can benefit the MTL performance (as elaborated in Appendix D.3), unsharing the classification head is the most effective approach, which leads to a desired behavior closely similar to the collaborative version of the linear models in Figure 1(a). As an additional experiment, we present Figure 4, which provides a similar comparison using an unfrozen version of ResNet-18 training from random initialization. Although this backbone is much deeper compared to the linear models we studied in Section 3.2, the consistent behavior suggests that our findings also generalize to more complex architectures.
@@
Figure 5: Test loss of the single-task vs multi-task learners on different datasets with pretrained backbones. The vertical axis is the cross-entropy loss and the horizontal axis is the width of the MLP on top of the backbone. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels. The multi-head architecture was used in all MTL experiments. These plots highlight the difference between MTL and STL double descent across several practical backbones and datasets.
@@
Additionally, Figure 5 offers further experiments on CUB and IN-R datasets trained on two different pretrained backbones. Observing similar phenomena to what we presented in observations 1-3, endorses our hypothesis regarding the connection between our results for linear models and DNNs in an MTL setup. For example, in all of these experiments the test error peak of the multi-task learner is considerably strong, while in some configurations, the STL error peak is negligible. This observation emphasizes that the label noise is not the sole source of the error peak; rather, as predicted from our theoretical analysis, the inherent dissimilarity across different tasks will also contribute to the strength of the peak. We conclude that our results help explain the behavior of DNNs, and combined with empirical investigation, they can be used for optimal design. Complementary experiments, as well as the full training and testing curves of MTL experiments, are provided in Appendix D.1.
@@
Figure 6 presents the test loss and the forgetting of a DNN continual learner on CIFAR-100 for different amounts of memory budget and model size. These models follow a trend very similar to the linear models presented in Section 4. In particular, there exists a test error peak for all models, but its location and strength are controlled by the memory size. By increasing the memory size, the test error peak naturally shifts to the right with reduced intensity. Another observation aligning with the linear models is that the forgetting greatly reduces for large model sizes, even without any memory, a phenomenon also confirmed by recent studies on sequential fine-tuning CL (Goldfarb & Hand, 2023).
Additionally, it is noticeable that increasing the memory capacity does not always uniformly increase the performance for all values of model size. Especially, the memory size is less effective with larger models. This raises a fundamental question when designing CL models in practice. Given a fixed budget for physical memory, is it more efficient to spend it on increasing the model size or building a bigger sample memory? Table 1 contains the results of experiments with varying budgets dedicated to the replay memory buffer. As the results show, different optimal balance points exist for all datasets that maximize the average accuracy and minimize forgetting. This observation aligns with the recent CL methods, where an equal portion of the sample memory budget is devoted to specialized blocks (Zhou et al., 2023; Wang et al., 2022) or task-specific tokens (Douillard et al., 2022). Therefore, we propose that our theoretical findings in linear models can be applied to practical CL algorithms to gain a deeper understanding.
@@
Figure 6: The test loss and forgetting of a continual learner on CIFAR-100 w.r.t. width of the model, $k$, for different levels of per-class memory capacity. $m=0$ means naive fine-tuning, and $m=500$ corresponds to the full MTL setup. (a) and (b) correspond to the experiments with the Pretrained ResNet-18 and subfigures (c) and (d) show the results when fine-tuning a ResNet-18 model from scratch. The multi-head architecture with no label noise was used in these experiments. These figures show that a CL learner also has a test error peak which is affected by the size of the memory buffer.
Table 1: The effect of varying sample memory budget and architecture memory budget over the average accuracy and average forgetting of overparameterized continual learners on two datasets. All values in this table are presented as percentages. The sample budget is linearly proportional to the number of training images stored in the memory, where 100% corresponds to storing 10% of the whole training set. A complete version of this table with more details is presented in Appendix C
. Mem. Budg. Arch. Budg. CIFAR-100 Imagenet-R CUB PMNIST Avg. Acc. Avg. Forg. Avg. Acc. Avg. Forg. Avg. Acc. Avg. Forg. Avg. Acc. Avg. Forg. 100 <1 $56.6_{\pm 0.4}$ $-22.5_{\pm 0.5}$ $76.1_{\pm 0.4}$ $-3.3_{\pm 0.6}$ $87.6_{\pm 0.2}$ $-4.5_{\pm 0.2}$ $32.8_{\pm 2.8}$ $-69.8_{\pm 3.0}$ 80 20 $70.6_{\pm 0.1}$ $-7.1_{\pm 0.1}$ $76.9_{\pm 0.6}$ $-2.0_{\pm 0.3}$ $84.1_{\pm 0.5}$ $-6.7_{\pm 0.0}$ $82.9_{\pm 1.0}$ $-13.9_{\pm 1.1}$ 60 40 $\bm{71.9_{\pm 0.2}}$ $\bm{-5.4_{\pm 0.3}}$ $77.1_{\pm 0.6}$ $-1.4_{\pm 0.4}$ $88.5_{\pm 0.3}$ $-1.8_{\pm 0.6}$ $87.7_{\pm 0.4}$ $-8.5_{\pm 0.4}$ 40 60 $66.3_{\pm 0.4}$ $-11.3_{\pm 0.4}$ $\bm{78.9_{\pm 0.5}}$ $-1.0_{\pm 0.1}$ $\bm{92.0_{\pm 0.1}}$ $\bm{-0.1_{\pm 0.2}}$ $\bm{88.8_{\pm 0.1}}$ $\bm{-7.1_{\pm 0.1}}$ 20 80 $67.4_{\pm 0.3}$ $-9.7_{\pm 0.3}$ $78.5_{\pm 0.3}$ $\bm{-0.6_{\pm 0.1}}$ $91.7_{\pm 0.4}$ $-0.4_{\pm 0.3}$ $88.2_{\pm 0.3}$ $-7.2_{\pm 0.2}$ 0 100 $63.8_{\pm 0.3}$ $-10.8_{\pm 0.3}$ $49.5_{\pm 0.6}$ $-20.5_{\pm 0.5}$ $75.8_{\pm 0.7}$ $-12.2_{\pm 0.7}$ $68.2_{\pm 1.0}$ $-17.9_{\pm 0.2}$